Find the mass of the following gas

175 mL CO2 at 0.711 atm and -3 degrees C


1
Expert's answer
2021-04-23T07:31:37-0400

"Mass= Unknown"

"Volume=175ml" "=0.175L"

"Pressure=0.711atm"

"Temperature=276k"

"0.08206L.atm\/K.mol"

From the equation;

"Pv=nRT"

"n=" "Pv\over RT"

"=" "(0.711\times 0.175)\over (0.08206 L.atm/K.mol)"

"=5.593\times 10^{-3}"

Mass to grams

"=5.595\times 10^{-3}mol\times 12.0g/mol"

"=0.067g"